OpenStudy (anonymous):

Help factor 2x^2+5x-168

OpenStudy (anonymous):

(2x+21)(x-8)

OpenStudy (aum):

2x^2+5x-168 = 2x^2 + 21x - 16x - 168 = x(2x+21) - 8(2x+21) = (2x+21)(x-8)